a) What is the maximum height you can suck water out of a open cup through a straw?

b) Instead of sucking through a long straw you instead fill a 2 L bottle with 1 L of water. You then seal a long straw to the bottle opening (i.e. only way in or out of the bottle is through the straw). If you blow one deep breath into the bottle, how high can the water raise through the straw? (Assume lung volume of 6 L and straw diameter of 1 cm.)
Answer #1

a) The maximum height one can suck out in an open cup is 10 m. This is because the water is pushed by air pressure and hence the force due gravity on the water column in the pipe can't be higher than the atmospheric pressure against the base of the water column.

(b) By using Boyle's law,
P1V1 = P2V2
P2 = P1*6
Let us assume P1 = atmospheric pressure = 101.3 kPa
Then
P2 = 607.8 kPa

P = pgh
=> h = P/pg
=> h = 607.8/(1000*9.8)
= 0.062 m